This is a question about "word problems involving quadratics." please help soon! The problem is, "the width of a flag is 3/5 it's length. The area of the flag is 1,500 square feet. What are the dimensions of the flag?" I am confused with how this problem is quadratic , and how to find the solution. I have attempted this problem multiple times! Thanks for the help.
Answer by richwmiller(17219)

The name "quadratic" comes from "quad" meaning square.
The length is squared.
We have the length L
and the width which is 3/4L
the flag's area is 1500
W*L=1500
3/4L*L=1500
divide by 3
L^2/4=500
L^2=2000
L=44.72